# Efficient Computation of Closed Substrings

### v1

@@ 5 MCSs in a Fibonacci Word
We prove the Lemma by induction on $n$. By direct counting for the base cases, we verify that $GM(f_{5})=1$ and $GM(f_{6})=2$. For the inductive step, assume that Equation (7) holds for $k\geq 7$. For $n\geq 3$, $f_{3}=101$ is a prefix of every $f_{n}$, and by Theorem 5.1 it is the only gapped-MCS in $f_{n}$. In $f_{k+1}=f_{k}f_{k-1}$, any new occurrences of $101$ in $f_{k+1}$ can only occur beginning in $f_{k}$ and ending in $f_{k-1}$. Moreover, an occurrence of $101$ as a suffix of $f_{k}$ or a prefix of $f_{k-1}$ may be lost by extension. Below we examine all such cases.
Suppose $k+1$ is odd. By inductive hypothesis, $GM(f_{k})=F_{k-5}+1$ and $GM(f_{k-1})=F_{k-6}$. Moreover, $f_{k}$ ends with the suffix $f_{2}=10$ and $f_{k-1}$ begins with $f_{3}=101$, resulting in the formation of the string $101$ crossing the boundary. Clearly, this string is not a gapped-MCS as it is right-extendible. The gapped-MCS $101$ which is the prefix of $f_{k-1}$ is no longer a gapped-MCS since $f_{k}$ ends with a 00 making it left-extendible. Thus, we get $GM(f_{k+1})=GM(f_{k})+GM(f_{k-1})-1=(F_{k-5}+1)+F_{k-6}-1=F_{k-4}$.
@@
Adding Equations (6),  (7) and the Equation in Lemma 6 and simplifying, we get Equation 8. Next, we have $M(f_{n})<F_{n}+F_{n-2}=\left(1+\frac{F_{n-2}}{F_{n}}\right)F_{n}=\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$ (since $\lim_{n\to\infty}\frac{F_{n-2}}{F_{n}}=\frac{1}{\phi^{2}}$). ∎
@@ 6 Experimental Analysis
All experiments were carried out on a server with an Intel Xeon Gold 6426Y CPU (64 cores, 128 threads, 75 MiB L3 cache), 250 GiB RAM, running RHEL 9.5 (kernel version 5.14.0-503.26.1). All code was implemented in C++ and compiled using GCC 11.5.0. The implementations for computing the suffix array ($\mathsf{SA}$) and the longest common prefix array ($\mathsf{LCP}$) were sourced from [18] and [21], respectively. The implementation is available at https://github.com/neerjamhaskar/closedStrings.
For $\sigma=2,3$ and $4$, we generate all strings of length $n$ for $1\leq n\leq 20$ and compute their MCSs. The results are plotted in Figure 2 in the Appendix. As observed, the maximum number of MCSs increases with the size of the alphabet $\sigma$ for strings of the same length. In total, the number of strings analyzed is equal to $\sum_{i=1}^{20}(2^{i}+3^{i}+4^{i})\approx 1.47$ trillion.
Acknowledgments We thank Simon J. Puglisi for introducing us to the MCS problem using $\mathsf{SA}$ and $\mathsf{LCP}$, which led to Theorem 4.4.
@@
$(2,4,7)$ $issi$, $issis$, $ississ$, $ississi$ $ississi$ Yes
@@
$(3,5,6)$ $ssiss$, $ssissi$ $ssissi$ No
$(3,1,2)$ $s$, $ss$ $ss$ Yes
$(4,5,5)$ $sissi$ $sissi$ No
$(4,3,3)$ $sis$ $sis$ Yes
$(4,1,1)$ $s$ $s$ Yes
$(5,4,4)$ $issi$ $issi$ No
@@
$(6,1,2)$ $s$, $ss$ $ss$ Yes
$(7,1,1)$ $s$ $s$ Yes
$(8,4,4)$ $ippi$ $ippi$ Yes

### v2

@@ 5 MCSs in a Fibonacci Word
We prove the Lemma by induction on $n$. By direct counting for the base cases, we verify that $GM(f_{5})=1$ and $GM(f_{6})=2$. For the inductive step, assume that Equation (7) holds for $k\geq 7$. For $n\geq 3$, $f_{3}=101$ is a prefix of every $f_{n}$, and by Theorem 5.1 every gapped-MCS in a Fibonacci word $f_{n}$, for $n\geq 5$, corresponds to an occurrence of the substring $101$ that is also a maximal gapped repeat. In $f_{k+1}=f_{k}f_{k-1}$, any new occurrences of $101$ in $f_{k+1}$ can only occur beginning in $f_{k}$ and ending in $f_{k-1}$. Moreover, an occurrence of $101$ as a suffix of $f_{k}$ or a prefix of $f_{k-1}$ may be lost by extension. Below we examine all such cases.
Suppose $k+1$ is odd. By inductive hypothesis, $GM(f_{k})=F_{k-5}+1$ and $GM(f_{k-1})=F_{k-6}$. Moreover, $f_{k}$ ends with the suffix $f_{2}=10$ and $f_{k-1}$ begins with $f_{3}=101$, resulting in the formation of the string $101$ crossing the boundary. Clearly, this string is not a gapped-MCS as it is right-extendible. The gapped-MCS $101$ which is the prefix of $f_{k-1}$ is no longer a gapped-MCS since $f_{k}$ ends with a 0 making it left-extendible. Thus, we get $GM(f_{k+1})=GM(f_{k})+GM(f_{k-1})-1=(F_{k-5}+1)+F_{k-6}-1=F_{k-4}$.
@@
Adding Equations (6),  (7) and the Equation in Lemma 6 and simplifying, we get Equation (8). Next, we have $M(f_{n})<F_{n}+F_{n-2}=\left(1+\frac{F_{n-2}}{F_{n}}\right)F_{n}\approx\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$ (since $\lim_{n\to\infty}\frac{F_{n-2}}{F_{n}}=\frac{1}{\phi^{2}}$). ∎
@@ 6 Experimental Analysis
All experiments were carried out on a server with an Intel Xeon Gold 6426Y CPU (64 cores, 128 threads, 75 MiB L3 cache), 250 GiB RAM, running RHEL 9.5 (kernel version 5.14.0-503.26.1). The code was implemented in C++ and compiled using GCC 11.5.0. The implementations for computing the suffix array ($\mathsf{SA}$) and the longest common prefix array ($\mathsf{LCP}$) were sourced from the libsais library [13], which is based on contributions from [15, 22, 25, 26]. The implementation is available at https://github.com/neerjamhaskar/closedStrings.
For $\sigma=2,3$ and $4$, we generate all strings of length $n$ for $1\leq n\leq 20$ and compute their MCSs. The results are plotted in Figure 2 in the Appendix. As observed, the maximum number of MCSs increases with the size of the alphabet $\sigma$ for strings of the same length. In total, the number of strings analyzed is equal to $\sum_{i=1}^{20}(2^{i}+3^{i}+4^{i})\approx 1.47$ trillion. These exhaustive computations raise a natural question: Given $n$ and $\sigma$, can we algorithmically construct a string that has the maximum possible number of MCSs?
{credits}
6.0.1 Acknowledgements
We thank Simon J. Puglisi for introducing us to the MCS problem using $\mathsf{SA}$ and $\mathsf{LCP}$, which led to Theorem 4.3. We are grateful to the reviewers for their valuable feedback.
@@ 99    $(2,1)$
$(2,4,7)$ issiissi, issisissis, ississississ, ississiississi ississiississi Yes
@@
$(3,5,6)$ ssissssiss, ssissississi ssissississi No
$(3,1,2)$ $s$, ssss ssss Yes
$(4,5,5)$ sissisissi sissisissi No
$(4,3,3)$ sissis sissis Yes
$(4,1,1)$ $s$ $s$ No
$(5,4,4)$ issiissi issiissi No
@@
$(6,1,2)$ $s$, ssss ssss Yes
$(7,1,1)$ $s$ $s$ No
$(8,4,4)$ ippiippi ippiippi Yes